Vector B~ has x, y, and z components of 6, 9.1, and 5.7 units, respectively. Calculate the magnitude of B~ .


1
Expert's answer
2021-10-04T10:06:26-0400

"\text{Magnitude of B}=\sqrt{(x-component)^2+(y-component)^2+(z-component)^2}\\
=\sqrt{6^2+9.1^2+5.7^2}=\text{12.3 units}"